# Mamba-Driven Topology Fusion for Monocular 3-D Human Pose Estimation

### v1

@@ Abstract
Transformer-based methods for 3-D human pose estimation face significant computational challenges due to the quadratic growth of self-attention mechanism complexity with sequence length. Recently, the Mamba model has substantially reduced computational overhead and demonstrated outstanding performance in modeling long sequences by leveraging state space model (SSM). However, the ability of SSM to process sequential data is not suitable for 3-D joint sequences with topological structures, and the causal convolution structure in Mamba also lacks insight into local joint relationships. To address these issues, we propose the Mamba-Driven Topology Fusion framework in this paper. Specifically, the proposed Bone Aware Module infers the direction and length of bone vectors in the spherical coordinate system, providing effective topological guidance for the Mamba model in processing joint sequences. Furthermore, we enhance the convolutional structure within the Mamba model by integrating forward and backward graph convolutional network, enabling it to better capture local joint dependencies. Finally, we design a Spatiotemporal Refinement Module to model both temporal and spatial relationships within the sequence. Through the incorporation of skeletal topology, our approach effectively alleviates Mamba’s limitations in capturing human structural relationships. We conduct extensive experiments on the Human3.6M and MPI-INF-3DHP datasets for testing and comparison, and the results show that the proposed method greatly reduces computational cost while achieving higher accuracy. Ablation studies further demonstrate the effectiveness of each proposed module. The code and models will be released.
Index Terms:
Monocular 3-D Human Pose Estimation, State Space Model, Human Topology Analysis, Graph Convolutional Network
@@ I Introduction
Monocular 3-D human pose estimation has garnered significant attention in the field of computer vision in recent years. The primary objective of this task is to estimate the 3-D coordinates of body joints from a given 2-D image or video captured from a single perspective. Research in this domain plays a pivotal role in enabling various downstream applications, such as action recognition [1], human-computer interaction [2], human body reconstruction [3, 4], and autonomous driving [5].
@@
Methods for single-view 3-D human pose estimation can be broadly classified into two categories: end-to-end methods and two-stage methods. End-to-end methods [6, 7] aim to directly recover 3-D joint coordinates from RGB images. In contrast, two-stage methods [8, 9, 10, 11] first estimate 2-D joint coordinates from RGB images and then use a lifting network to map the 2-D pose to its 3-D counterpart. Given the outstanding performance of 2-D pose detectors [12, 13, 14], two-stage methods often achieve higher accuracy in 3-D pose estimation, making them the primary focus of this discussion. However, transforming 2-D poses into 3-D coordinates presents significant challenges, such as the loss of depth information and issues related to self-occlusion. To address these challenges, convolutional neural network (CNN) [15, 16, 17] have been widely used to extract depth features from images. More recently, due to the exceptional success of transformer [18] in computer vision, transformer-based architectures have been extensively adopted in 3-D human pose estimation and demonstrated superior performance compared to CNN-based models. For example, PoseFormer [8] was the first to introduce transformer into 3-D human pose estimation, exploring both spatial and temporal information in pose sequences. MotionBERT [19] designed a dual-stream transformer architecture and achieved SOTA performance by incorporating pre-training techniques. HCT [20] utilized hierarchical cross transformer to encode high-level temporal, spatial and cross-view information to improve unsupervised 3-D HPE. Although these methods achieved promising results, they introduced significant computational and memory overhead due to the quadratic complexity of the self-attention mechanism with respect to the length of pose sequences. Approaches such as dividing time frames [21, 22] and removing redundant sequences [23, 24, 25] can alleviate this issue but fail to fundamentally solve the problem.
@@
The structured state space model (SSM) [26] has recently demonstrated linear scaling with respect to sequence length and achieved state-of-the-art (SOTA) performance across several modalities, including language, speech, and genomics. Mamba [27] is an improvement over SSM, incorporating an input-dependent selection mechanism and an efficient hardware-aware design. Given the Mamba model’s outstanding capabilities in long-sequence modeling, its application to 3-D pose estimation tasks, which heavily rely on long pose temporal sequences, is highly effective. Several existing methods that incorporate the Mamba into 3-D HPE, such as PoseMamba [28], PoseMagic [29], and HGMamba [30], have achieved competitive performance. However, directly applying the Mamba to capture the spatial structure of human posture is not appropriate. In contrast to the self-attention mechanism, which simultaneously attends to all tokens globally, SSM processes sequential data relying only on tokens preceding the current position. However, the spatial configuration of human joints is inherently non-sequential, as it adheres to the intrinsic skeletal topology of the human body. Moreover, prior to applying SSM, Mamba employs causal convolutions to capture local dependencies within the sequence. This causal convolution strategy is also suboptimal for modeling the complex structural relationships inherent in the human body.
@@
In the spatiotemporal refinement module, drawing upon the encoder block architecture of Vision Mamba [31], we propose a upgraded module termed GCN-Enhanced Vision Mamba (GEM). Graph Convolutional Network (GCN) [32] have proven effective for modeling sequential data with inherent graph structures, and have been integrated with Mamba in several recent approaches, including PoseMagic [29], HGMamba [30], and Hamba [33]. Unlike these methods, where GCN and Mamba are implemented as independent components, we modify the internal convolutional architecture of the Mamba model and incorporate both forward and backward graph convolutional network. GEM effectively captures both global and local dependencies among joints and demonstrates superior performance compared to alternative approaches. Multiple GEM and Vision Mamba blocks are stacked to form the spatiotemporal refinement module. Furthermore, we propose a bone-joint fusion embedding mechanism to embed and integrate the features of 2-D joints and bone vectors. The overall architecture of our framework is illustrated in Figure 1.
@@
  • We propose a spatiotemporal refinement module for accurate 3-D pose estimation. Additionally, we enhance the convolutional architecture within the Mamba model by integrating forward and backward GCN, which facilitates the capture of local topological dependencies among skeletal joints.
@@ II Related Work
II-A Two-stage 3-D Human Pose Estimation
Two-stage 3-D human pose estimators typically extract 2-D pose information from images, followed by a 2-D to 3-D booster. Most methods focus on the design of the booster. Martinez et al. [36] used a multi-layer fully connected neural network to elevate 2-D coordinates to 3-D. Fang et al. [37] enhanced high-level constraints on human pose in 2-D to 3-D lifting using pose grammar. Wandt et al. [38] trained a 3-D to 2-D reprojection network using adversarial training. Li et al. [39] applied data augmentation and designed a deep cascaded fully connected neural network to regress 3D coordinates. Due to the powerful global modeling capability of transformers, they can significantly improve accuracy in 3-D HPE tasks. Shan et al. [9] employed pretraining and fine-tuning to better capture joints dependencies. Zhang et al. [11] proposed a spatiotemporal alternating network structure to capture joint spatial correlation and temporal motion. Some methods focus on exploring trajectory information across pose time frames. Pavllo et al. [40] used dilated temporal convolutions to explore frame relationships. Li et al. [23] introduced a strided transformer to predict single-frame poses from a sequence. Other methods [19, 41, 42, 43, 10] also explore temporal information within models to improve accuracy. However, while transformer-based methods enhance model performance, they also reduce execution efficiency. Although some works [44, 25, 24, 45] aim to reduce computational cost, the quadratic computational complexity of the attention mechanism remains a challenge.
@@ II-B State Space Model
Recently, state space model (SSM) have shown great potential in modeling long-range sequences across multiple modalities. The Structured State Space Sequence model (S4) [26] improves upon SSM by using HiPPO matrices with conditioning matrices $A$, incorporating discretization techniques, and convolutional representations. Mamba [27] further enhances S4 by introducing parameter-dependent improvements in SSM and employing a parallel scanning algorithm for efficient model computation. Subsequently, Vision Mamba [31] proposed a general-purpose visual backbone with bidirectional Mamba blocks, and VMamba [46] utilized 2D Selective Scan (SS2D) along four scanning paths to help understand the non-sequential structure of 2-D visual data. DVMSR [47] employed Vision Mamba and a distillation strategy to achieve state-of-the-art (SOTA) performance in image super-resolution tasks. PoseMamba [28] and PoseMagic [29] employ the Mamba model for 3-D HPE, achieving reduced computational cost while maintaining favorable performance. Motivated by the strong performance of Mamba, we further explore its topological capabilities in the context of 3-D HPE.
II-C Human Topology Analysis in 3-D HPE
In 3-D HPE, analyzing the human topology and processing different structures to reduce errors is a common approach. Xu et al. [48] decoupled the 3-D pose into bone lengths and bone orientations, predicting them separately using a two-stream structure. Cai et al. [49] employed a diffusion model with Transformers for denoising noisy bone lengths and orientations, targeting parent and child joints. Xue et al. [50] divided the human body into five parts, exploring their temporal dependencies and relationships with other actions. Furthermore, Graph Convolutional Networks (GCN) effectively represent the topological relationships between human joints using adjacency matrices. Yin et al. [20] utilized multi-branch attention graph convolution to extract feature information contributing to different nodes. Cai et al. [51] utilized GCN-based methods to integrate domain knowledge of human body layout. While GCN is a lightweight model, it often lacks high accuracy. Wang et al. [52] proposed an advanced method for handling joints with different degrees of freedom, but it cannot impose strong constraints on all joints. Wu et al. [22] applied relative and absolute bone angle constraints on limbs; however, this limitation is flawed as identical relative and absolute bone angles may occur in different actions. Inspired by these methods, we propose using bone vectors in spherical coordinates as global topological information for the human body, where the relationship between two joints is both intuitive and unique.
@@ III Method
Figure 1: The overall structure of the Mamba-Driven Topology Fusion framework. We perform seq2seq estimation using 2-D pose sequences as input. $s_{2D}$ represents the 2-D pose sequence detected from images. First, the bone aware module predicts the category of bone vectors based on $s_{2D}$, and the corresponding bone vector coordinates are then computed by combining the predicted categories with the 2-D joint positions. Then, the bone vectors and 2-D poses are combined and fed into the bone-joint fusion embedding. Finally, the spatiotemporal refinement module regresses accurate 3-D pose sequences.
The Mamba-Driven Topology Fusion framework we propose is shown in Figure 1. Our input is a 2-D pose sequence frame $s_{2D}$ obtained from the 2-D pose detector. We first train the bone aware module to compute the polar angle categories for each bone in the spherical coordinate system. Then, we assign values to these categories and infer the estimated spherical coordinates $bone_{s}\in\mathbb{R}^{f\times j\times 3}$ for each bone based on the human body structure, where the last dimension corresponds to $r$, representing the radial distance, $\theta$, representing the polar angle, and $\phi$, representing the azimuthal angle. Next, we feed $s_{2D}$ and $bone_{s}$ into the bone-joint fusion embedding for feature fusion, followed by the spatiotemporal refinement module to capture both spatial and temporal information of the joints. Finally, a regression head is used to predict the true 3-D pose $s_{3D}\in\mathbb{R}^{f\times j\times 3}$. In the following section, we first present the preliminaries of SSM, followed by a description of the Mamba-based GEM block, and then detail the proposed components.
@@ III-B GCN-Enhanced Vision Mamba
Figure 2: The detailed components of the VIM and GEM blocks are illustrated. Subfigure a illustrates the overall architecture of VIM along with its internal $Vim$ module. Subfigure b presents the structure of GEM and details the convolutional components enhanced with graph convolution.
@@
$\tilde{x}_{f}$ $=SSM_{f}(\sigma(Conv1d_{f}(\hat{x})))$ (4)
$\tilde{x}_{b}$ $=SSM_{b}(\sigma(Conv1d_{b}(flip(\hat{x}))))$
@@
$y_{f}$ $=\tilde{x}_{f}\circ SiLU(\hat{z})$ (5)
$y_{b}$ $=flip(\tilde{x}_{b})\circ SiLU(\hat{z})$
@@
$\bar{x}_{b}=SSM_{b}(\sigma(Conv1d_{b}(GCN_{b}(LN(flip(\hat{x}))))))$ (9)
@@ III-C Bone Aware Module
The bone aware module computes the direction and length of bone vectors, providing human skeletal topology as guidance for Mamba during joint position inference. This partially addresses Mamba’s limited capacity to capture human body topology when processing the entire pose sequence. The bone vector in 3-D space contains depth positioning information of the child joint relative to the parent joint. As shown in Figure 3, a bone vector in space is defined as $l_{i}=p^{child}_{i}-p^{parent}_{i}$, where $p^{child}_{i}$ and $p^{parent}_{i}$ represent the child and parent joints of the $i$-th bone, respectively. This vector is then transferred to the spherical coordinate system, where the parent joint $p^{parent}_{i}$ is aligned with the origin $o$ of the spherical coordinate system, and the child joint can be represented by spherical coordinates $(r_{i},\theta_{i},\phi_{i})$. If the spherical coordinates of the child joint for each bone vector can be estimated, the direction and length of the bone vector can be roughly determined, which helps guide the model to capture the correct human body topology.
@@
where $z_{i}$ is unknown, $x_{i}$ and $y_{i}$ can be obtained from the 2-D pose information, with $\theta_{i}$ ranging from 0 to $\pi$ and $\phi_{i}$ ranging from 0 to $2\pi$. In Equation 11, the polar angle $\theta_{i}$ contains depth information, and inferring the polar angle allows further calculation of the radial distance $r_{i}$. However, directly regressing the polar angle is challenging, so we perform classification to predict it. By dividing the range of the polar angle into categories of size $\pi/n$, we can classify it into $n$ categories, each representing an approximate direction of the bone’s depth.
Next, we design the bone aware module to predict the polar angle category of each bone vector in the human skeleton. It takes a 2-D pose sequence $s_{2D}$ as input, followed by a pyramid-shaped classification network with feature attenuation. As shown in Figure 3, given $s_{2D}$, we first apply a feature embedding to map it to a high-dimensional feature $s^{\prime}_{2D}\in\mathbb{R}^{f\times j\times d}$. The network of the bone aware module consists of $l$ layers, each containing a spatial VIM, a temporal VIM, and a feature decay block. For the computation at the $i$-th layer, we first use VIM to capture the spatial features of the bone structure:
@@
where $s^{t}_{i}\in\mathbb{R}^{j\times f\times d}$, and $trans$ represents the transpose operation applied to the temporal and spatial dimensions. To reduce the computational load and compress the model features, we apply a feature attenuation matrix to $s^{t}_{i}$:

### v2

# Mamba-Driven Topology Fusion for Monocular 3D Human Pose Estimation

@@ Abstract
Transformer-based methods for 3D human pose estimation face significant computational challenges due to the quadratic growth of self-attention mechanism complexity with sequence length. Recently, the Mamba model has substantially reduced computational overhead and demonstrated outstanding performance in modeling long sequences by leveraging state space model (SSM). However, the ability of SSM to process sequential data is not suitable for 3D joint sequences with topological structures, and the causal convolution structure in Mamba also lacks insight into local joint relationships. To address these issues, we propose the Mamba-Driven Topology Fusion framework in this paper. Specifically, the proposed Bone Aware Module infers the direction and length of bone vectors in the spherical coordinate system, providing effective topological guidance for the Mamba model in processing joint sequences. Furthermore, we enhance the convolutional structure within the Mamba model by integrating forward and backward graph convolutional network, enabling it to better capture local joint dependencies. Finally, we design a Spatiotemporal Refinement Module to model both temporal and spatial relationships within the sequence. Through the incorporation of skeletal topology, our approach effectively alleviates Mamba’s limitations in capturing human structural relationships. We conduct extensive experiments on the Human3.6M and MPI-INF-3DHP datasets for testing and comparison, and the results show that the proposed method greatly reduces computational cost while achieving higher accuracy. Ablation studies further demonstrate the effectiveness of each proposed module. The code and models will be released.
@@ I Introduction
Monocular 3D human pose estimation has garnered significant attention in the field of computer vision in recent years. The primary objective of this task is to estimate the 3D coordinates of body joints from a given 2D image or video captured from a single perspective. Research in this domain plays a pivotal role in enabling various downstream applications, such as action recognition [1], human-computer interaction [2], human body reconstruction [3, 4], and autonomous driving [5].
@@
Methods for single-view 3D human pose estimation can be broadly classified into two categories: end-to-end methods and two-stage methods. End-to-end methods [6, 7] aim to directly recover 3D joint coordinates from RGB images. In contrast, two-stage methods [8, 9, 10, 11] first estimate 2D joint coordinates from RGB images and then use a lifting network to map the 2D pose to its 3D counterpart. Given the outstanding performance of 2D pose detectors [12, 13, 14], two-stage methods often achieve higher accuracy in 3D pose estimation, making them the primary focus of this discussion. However, transforming 2D poses into 3D coordinates presents significant challenges, such as the loss of depth information and issues related to self-occlusion. To address these challenges, convolutional neural network (CNN) [15, 16, 17] have been widely used to extract depth features from images. More recently, due to the exceptional success of transformer [18] in computer vision, transformer-based architectures have been extensively adopted in 3D human pose estimation and demonstrated superior performance compared to CNN-based models. For example, PoseFormer [8] was the first to introduce transformer into 3D human pose estimation, exploring both spatial and temporal information in pose sequences. MotionBERT [19] designed a dual-stream transformer architecture and achieved SOTA performance by incorporating pre-training techniques. HCT [20] utilized hierarchical cross transformer to encode high-level temporal, spatial and cross-view information to improve unsupervised 3D HPE. Although these methods achieved promising results, they introduced significant computational and memory overhead due to the quadratic complexity of the self-attention mechanism with respect to the length of pose sequences. Approaches such as dividing time frames [21, 22] and removing redundant sequences [23, 24, 25] can alleviate this issue but fail to fundamentally solve the problem.
@@
The structured state space model (SSM) [26] has recently demonstrated linear scaling with respect to sequence length and achieved state-of-the-art (SOTA) performance across several modalities, including language, speech, and genomics. Mamba [27] is an improvement over SSM, incorporating an input-dependent selection mechanism and an efficient hardware-aware design. Given the Mamba model’s outstanding capabilities in long-sequence modeling, its application to 3D pose estimation tasks, which heavily rely on long pose temporal sequences, is highly effective. Several existing methods that incorporate the Mamba into 3D HPE, such as PoseMamba [28], PoseMagic [29], and HGMamba [30], have achieved competitive performance. However, directly applying the Mamba to capture the spatial structure of human posture is not appropriate. In contrast to the self-attention mechanism, which simultaneously attends to all tokens globally, SSM processes sequential data relying only on tokens preceding the current position. However, the spatial configuration of human joints is inherently non-sequential, as it adheres to the intrinsic skeletal topology of the human body. Moreover, prior to applying SSM, Mamba employs causal convolutions to capture local dependencies within the sequence. This causal convolution strategy is also suboptimal for modeling the complex structural relationships inherent in the human body.
@@
In the spatiotemporal refinement module, drawing upon the encoder block architecture of Vision Mamba [31], we propose a upgraded module termed GCN-Enhanced Vision Mamba (GEM). Graph Convolutional Network (GCN) [32] have proven effective for modeling sequential data with inherent graph structures, and have been integrated with Mamba in several recent approaches, including PoseMagic [29], HGMamba [30], and Hamba [33]. Unlike these methods, where GCN and Mamba are implemented as independent components, we modify the internal convolutional architecture of the Mamba model and incorporate both forward and backward graph convolutional network. GEM effectively captures both global and local dependencies among joints and demonstrates superior performance compared to alternative approaches. Multiple GEM and Vision Mamba blocks are stacked to form the spatiotemporal refinement module. Furthermore, we propose a bone-joint fusion embedding mechanism to embed and integrate the features of 2D joints and bone vectors. The overall architecture of our framework is illustrated in Figure 1.
@@
  • We propose a spatiotemporal refinement module for accurate 3D pose estimation. Additionally, we enhance the convolutional architecture within the Mamba model by integrating forward and backward GCN, which facilitates the capture of local topological dependencies among skeletal joints.
@@ II Related Work
II-A Two-stage 3D Human Pose Estimation
Two-stage 3D human pose estimators typically extract 2D pose information from images, followed by a 2D to 3D booster. Most methods focus on the design of the booster. Martinez et al. [36] used a multi-layer fully connected neural network to elevate 2D coordinates to 3D. Fang et al. [37] enhanced high-level constraints on human pose in 2D to 3D lifting using pose grammar. Wandt et al. [38] trained a 3D to 2D reprojection network using adversarial training. Li et al. [39] applied data augmentation and designed a deep cascaded fully connected neural network to regress 3D coordinates. Due to the powerful global modeling capability of transformers, they can significantly improve accuracy in 3D HPE tasks. Shan et al. [9] employed pretraining and fine-tuning to better capture joints dependencies. Zhang et al. [11] proposed a spatiotemporal alternating network structure to capture joint spatial correlation and temporal motion. Some methods focus on exploring trajectory information across pose time frames. Pavllo et al. [40] used dilated temporal convolutions to explore frame relationships. Li et al. [23] introduced a strided transformer to predict single-frame poses from a sequence. Other methods [19, 41, 42, 43, 10] also explore temporal information within models to improve accuracy. However, while transformer-based methods enhance model performance, they also reduce execution efficiency. Although some works [44, 25, 24, 45] aim to reduce computational cost, the quadratic computational complexity of the attention mechanism remains a challenge.
@@ II-B State Space Model
Recently, state space model (SSM) have shown great potential in modeling long-range sequences across multiple modalities. The Structured State Space Sequence model (S4) [26] improves upon SSM by using HiPPO matrices with conditioning matrices $A$, incorporating discretization techniques, and convolutional representations. Mamba [27] further enhances S4 by introducing parameter-dependent improvements in SSM and employing a parallel scanning algorithm for efficient model computation. Subsequently, Vision Mamba [31] proposed a general-purpose visual backbone with bidirectional Mamba blocks, and VMamba [46] utilized 2D Selective Scan (SS2D) along four scanning paths to help understand the non-sequential structure of 2D visual data. DVMSR [47] employed Vision Mamba and a distillation strategy to achieve state-of-the-art (SOTA) performance in image super-resolution tasks. PoseMamba [28] and PoseMagic [29] employ the Mamba model for 3D HPE, achieving reduced computational cost while maintaining favorable performance. Motivated by the strong performance of Mamba, we further explore its topological capabilities in the context of 3D HPE.
II-C Human Topology Analysis in 3D HPE
In 3D HPE, analyzing the human topology and processing different structures to reduce errors is a common approach. Xu et al. [48] decoupled the 3D pose into bone lengths and bone orientations, predicting them separately using a two-stream structure. Cai et al. [49] employed a diffusion model with Transformers for denoising noisy bone lengths and orientations, targeting parent and child joints. Xue et al. [50] divided the human body into five parts, exploring their temporal dependencies and relationships with other actions. Furthermore, Graph Convolutional Networks (GCN) effectively represent the topological relationships between human joints using adjacency matrices. Yin et al. [20] utilized multi-branch attention graph convolution to extract feature information contributing to different nodes. Cai et al. [51] utilized GCN-based methods to integrate domain knowledge of human body layout. While GCN is a lightweight model, it often lacks high accuracy. Wang et al. [52] proposed an advanced method for handling joints with different degrees of freedom, but it cannot impose strong constraints on all joints. Wu et al. [22] applied relative and absolute bone angle constraints on limbs; however, this limitation is flawed as identical relative and absolute bone angles may occur in different actions. Inspired by these methods, we propose using bone vectors in spherical coordinates as global topological information for the human body, where the relationship between two joints is both intuitive and unique.
@@ III Method
Figure 1: The overall structure of the Mamba-Driven Topology Fusion framework. We perform seq2seq estimation using 2D pose sequences as input. $s_{2D}$ represents the 2D pose sequence detected from images. First, the bone aware module predicts the category of bone vectors based on $s_{2D}$, and the corresponding bone vector coordinates are then computed by combining the predicted categories with the 2D joint positions. Then, the bone vectors and 2D poses are combined and fed into the bone-joint fusion embedding. Finally, the spatiotemporal refinement module regresses accurate 3D pose sequences.
The Mamba-Driven Topology Fusion framework we propose is shown in Figure 1. Our input is a 2D pose sequence frame $s_{2D}$ obtained from the 2D pose detector. We first train the bone aware module to compute the polar angle categories for each bone in the spherical coordinate system. Then, we assign values to these categories and infer the estimated spherical coordinates $bone_{s}\in\mathbb{R}^{f\times j\times 3}$ for each bone based on the human body structure, where the last dimension corresponds to $r$, representing the radial distance, $\theta$, representing the polar angle, and $\phi$, representing the azimuthal angle. Next, we feed $s_{2D}$ and $bone_{s}$ into the bone-joint fusion embedding for feature fusion, followed by the spatiotemporal refinement module to capture both spatial and temporal information of the joints. Finally, a regression head is used to predict the true 3D pose $s_{3D}\in\mathbb{R}^{f\times j\times 3}$. In the following section, we first present the preliminaries of SSM, followed by a description of the Mamba-based GEM block, and then detail the proposed components.
@@ III-B GCN-Enhanced Vision Mamba
Figure 2: The detailed components of the VIM and GEM blocks are illustrated. Subfigure a illustrates the overall architecture of VIM along with its internal VimVim module. Subfigure b presents the structure of GEM and details the convolutional components enhanced with graph convolution.
@@
[…]
$\tilde{x}_{b}$ $))\displaystyle=SSM_{b}(\sigma(Conv1d_{b}(flip(\hat{x}$))))
@@
[…]
@@
$)))\bar{x}_{b}=SSM_{b}(\sigma(Conv1d_{b}(GCN_{b}(LN(flip(\hat{x}$)))))) (9)
@@ III-C Bone Aware Module
The bone aware module computes the direction and length of bone vectors, providing human skeletal topology as guidance for Mamba during joint position inference. This partially addresses Mamba’s limited capacity to capture human body topology when processing the entire pose sequence. The bone vector in 3D space contains depth positioning information of the child joint relative to the parent joint. As shown in Figure 3, a bone vector in space is defined as $l_{i}=p^{child}_{i}-p^{parent}_{i}$, where $p^{child}_{i}$ and $p^{parent}_{i}$ represent the child and parent joints of the $i$-th bone, respectively. This vector is then transferred to the spherical coordinate system, where the parent joint $p^{parent}_{i}$ is aligned with the origin $o$ of the spherical coordinate system, and the child joint can be represented by spherical coordinates $(r_{i},\theta_{i},\phi_{i})$. If the spherical coordinates of the child joint for each bone vector can be estimated, the direction and length of the bone vector can be roughly determined, which helps guide the model to capture the correct human body topology.
@@
where $z_{i}$ is unknown, $x_{i}$ and $y_{i}$ can be obtained from the 2D pose information, with $\theta_{i}$ ranging from 0 to $\pi$ and $\phi_{i}$ ranging from 0 to $2\pi$. In Equation 11, the polar angle $\theta_{i}$ contains depth information, and inferring the polar angle allows further calculation of the radial distance $r_{i}$. However, directly regressing the polar angle is challenging, so we perform classification to predict it. By dividing the range of the polar angle into categories of size $\pi/n$, we can classify it into $n$ categories, each representing an approximate direction of the bone’s depth.
Next, we design the bone aware module to predict the polar angle category of each bone vector in the human skeleton. It takes a 2D pose sequence $s_{2D}$ as input, followed by a pyramid-shaped classification network with feature attenuation. As shown in Figure 3, given $s_{2D}$, we first apply a feature embedding to map it to a high-dimensional feature $s^{\prime}_{2D}\in\mathbb{R}^{f\times j\times d}$. The network of the bone aware module consists of $l$ layers, each containing a spatial VIM, a temporal VIM, and a feature decay block. For the computation at the $i$-th layer, we first use VIM to capture the spatial features of the bone structure:
@@
where $s^{t}_{i}\in\mathbb{R}^{j\times f\times d}$, and transtrans represents the transpose operation applied to the temporal and spatial dimensions. To reduce the computational load and compress the model features, we apply a feature attenuation matrix to $s^{t}_{i}$: